OpenStudy (anonymous):

Choose the correct description of the graph of the inequality -x < 3 - Open circle on 3, shading to the left. Closed circle on 3, shading to the left. Open circle on -3, shading to the right. Closed circle on -3, shading to the right.

OpenStudy (anonymous):

Again, whaddya think?

OpenStudy (anonymous):

first one?

OpenStudy (anonymous):

First of all, \[\geq or\leq \]means you have to include that number (closed circle).

OpenStudy (anonymous):

so the 2nd one ?

OpenStudy (anonymous):

second of all, write \[-x\leq 3 \;\;as \;\;x\geq -3\] for convenience.

OpenStudy (anonymous):

ohh wait its the last one :)

OpenStudy (anonymous):

A hit on the third throw!!!